Two bulbs A and B are rated 100 W - 120 W and 10W - 120V respectively.they are connected across a 120V source in series.which will consume more energy?
1
Expert's answer
2012-07-17T09:31:41-0400
Let's find out the resistances of each bulb from the data given:
R1 = U1^2 /
P1 = 120^2 / 100 = 144 Ohm
R2 = U2^2 / P2 = 120^2 / 10 = 1440 Ohm
Let's
use P = I^2 * R for each bulb and find out if it is greater.
P1 = (120 / (144
+ 1440))^2 * 144 = 0.8264 W
P2 = (120 / (144 + 1440))^2 * 1440 = 8.2644
W.
Answer: second bulb will consume more energy.